# AquaFuse: Waterbody Fusion for Physics Guided View Synthesis of Underwater Scenes

### v1

@@ 1 Introduction
Light attenuates exponentially underwater with propagation distance [2, 44] due to two physical characteristics of light: scattering and absorption. Forward scattering is responsible for blur whereas backscatter causes contrast reduction [43, 17]. Besides, light absorption by water has a prominent spectral dependency, which depends on the specific optical properties of a waterbody [3], distance of light sources, salinity, and many other factors. These physical characteristics are represented by the underwater image formation model [2], which, in recent years, have been successfully applied for image restoration [3] and scene rendering tasks [20, 41].
In particular, underwater scene synthesis approaches extend the state-of-the-art (SOTA) neural rendering methods such as NeRF [30] and Gaussian splatting [19] for domain-aware learning. They incorporate the physical properties of underwater image formation into the rendering pipeline for accurate view synthesis [20, 41]. This helps compensate for the inherent attenuation and optical artifacts in underwater imagery and ensures geometrically consistency. Various image recognition models and 3D reconstruction pipelines also use them for improved visual perception [43, 38].
@@
Note that the image style transfer literature [8, 46] offers solutions that learn artistic or photometric features from data. These methods rely on generative or adversarial networks to adapt one image domain’s style to another [10, 26]. While they excel at perceptual enhancements and artistic blending [17], physical constraints of underwater image formation [2] are not preserved. Hence, data augmentation by such domain transfer is not suitable if physical realism and object geometry preservation are crucial, as in 3D view synthesis, scene reconstructions, robotic mapping, and visual servoing tasks [41, 39].
@@
In addition to the theoretical analysis, we present the engineering constructs to tackle the practicalities of waterbody fusion across multiple scenes. We demonstrate how AquaFuse can be used for waterbody crossover, image enhancement by fusing with a clear reference scene, and in a generative pipeline for data augmentation. For qualitative and quantitative validation, we collect data at multiple ocean sites over diverse waterbody types and depth levels ($10^{\prime}$-$80^{\prime}$). Our experiments show that AquaFuse (i) performs perceptually realistic waterbody fusion; (ii) ensures geometric consistency across underwater scenes; (iii) is invariant to dataset bias and overfitting unlike learning-based approaches; and (iv) offers fast end-to-end computational being a closed-form solution.
In particular, our analyses of $N$ images across six water body types show that AquaFused images preserve over $94\%$ depth consistency and $90$-$95\%$ structural similarity of input scenes. The performance margins vary $1$-$5$% with increasing water depths from $10^{\prime}$ to $80^{\prime}$. Moreover, we integrate AquaFuse within a 3D view synthesis pipeline using Gaussian Splatting [19] for underwater scene reconstruction across multiple waterbodies. We find that AquaFuse can generate highly detailed 3D reconstructions that accurately represent the original scene’s geometric characteristics in the fused scenes.
As annotated underwater image databases are scarce, physics-guided data augmentation by AquaFuse can facilitate a multi-fold increase in training samples, which can be used for both 2D image recognition and 3D scene reconstruction [41, 45, 20]. This covers a wide spectrum of robot vision tasks, empowering learning-based models to be more robust and generalizable [44, 1]. The computational efficiency of AquaFuse also enables it to run in real-time for active vision, which data-driven style transfer approaches generally fail to offer.
@@ 2 Background & Preliminaries
Here, $\mathbf{\lambda}$ denotes the different wavelengths for RGB channels and $d(\mathbf{x})$ is the distance of point $\mathbf{x}$ along the line-of-sight. The transmission map of water medium is an exponential of the form $t_{\mathbf{\lambda}}(\mathbf{x})=e^{-\beta_{\lambda}\cdot d(\mathbf{x})}$, with $\beta_{\mathbf{\lambda}}$ as the uniform attenuation coefficient. $\mathbf{B}_{\mathbf{\lambda}}^{\infty}$ is the background (veiling) light.
The revised underwater image formation model [2, 3] shows that the effects of light absorption can be comparable to light scattering or even dominate it depending on the waterbody type. Besides, the attenuation coefficient relies not only on the optical properties of water, but also on the sensor, scene irradiance, and imaging range. More importantly, attenuation coefficients of backscattering and direct signal are different, which can be reduced to $\beta_{\mathbf{\lambda}}^{B}=-\ln(1-{\mathbf{B}_{\mathbf{\lambda}}}/{\mathbf{B}_{\mathbf{\lambda}}^{\infty}})/\mathbf{z}$ and $\beta_{\mathbf{\lambda}}^{D}=-\ln({(\mathbf{U}_{\mathbf{\lambda}}-\mathbf{B}_{\mathbf{\lambda}})}/{\mathbf{I}_{\mathbf{\lambda}}})/\mathbf{z}$, respectively. The revised model is then expressed by:
@@
where $v_{D}=\{z,\rho,{E},S_{\lambda},\beta\}$ is the coefficient dependency of direct signal and $v_{B}=\{{E},S_{\lambda},b,\beta\}$ is coefficient dependency of backscattered signal. The parameters $\rho$ is the reflectance, ${E}$ is the spectrum of ambient light, $S_{\lambda}$ is spectral response, $b$ is beam scattering coefficient, and $\beta$ is beam attenuation coefficient.
In recent years, this revised image formation model has been successfully used for rendering underwater scenes by methods such as SeaThru-NeRF [20], WaterSplatting [25], and SeaSplat [41]. These methods leverage the remarkable advances in 3D radiance fields, specifically NeRF [30] and Gaussian splatting [19]-based rendering. However, drawbacks remain in accounting for multiple scattering or inconsistent/artificial illumination. They also require camera poses to be extracted beforehand which is often infeasible. Thus, collecting and synthesizing data for 3D underwater scene reconstruction research is rather challenging.
To this end, geometric and photometric data augmentations with various perspective warping generally help [45]. This paper introduces another photometric augmentation technique for waterbody fusion to improve the generalizability of learning underwater image recognition and reconstruction tasks.
@@ 3 AquaFuse Methodology
Figure 2: An example of our scene restoration pipeline is shown; (a) raw input image $\mathbf{U}$; (b) estimated depth $\mathbf{Z}$; (c) recovered image $\mathbf{I}$; and (d-e) illumination maps for R-G-B channels, respectively.
@@ 3.1 Real-time Depth Estimation
A sample example of estimated $\mathbf{Z}\sim\tilde{d}(\cdot)$ is shown in Fig. 2. In our implementation, an additional median blur is applied to reduce noise and preserve edges, which ensures the robustness of background light estimation (see Sec. 3.2). Specifically, we perform $\mathbf{z}=MedianBlur(\mathbf{z},7)$ to filter the estimated depth map with a kernel size of $7$.
@@ 4.3.4 Geometry-preserving Scene Transformation
These visual inspections corroborate our quantitative analyses of the reconstructed scenes. As shown in Table 2, the PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS) scores of AquaFused images are within $1$-$2$% of original scenes. In fact, AquaFuse slightly improves 3D reconstruction quality with a higher PSNR and lower LPIPS at shallow depths ($10^{\prime}$-$30^{\prime}$). The performance margins are slightly lower for deep-water scenes (over $60^{\prime}$); nevertheless, the scores are very similar to using the original images’ reconstruction quality, which validates our intuition and design objectives.
Table 2: 3D reconstruction (Gaussian Splatting) performance of AquaFused images versus original images at varied water depths.
Depth Image PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
$10^{\prime}$ Original $22.83$ $0.8104$ $0.2289$

### v2

@@ 1 Introduction
Light attenuates exponentially underwater with propagation distance [2, 48] due to two physical characteristics of light: scattering and absorption. Forward scattering is responsible for blur, whereas backscatter causes contrast reduction [47, 22]. Besides, light absorption by water has a prominent spectral dependency, which depends on the specific optical properties of a waterbody [3], distance of light sources, salinity, and many other factors. These physical characteristics are represented by the underwater image formation model [2], which, in recent years, has been successfully applied for water removal, image restoration [3], and various scene rendering tasks [25, 45]. In particular, underwater scene synthesis approaches extend the state-of-the-art (SOTA) neural rendering methods such as NeRF [35] and Gaussian splatting [24] for domain-aware learning. They incorporate the physical properties of underwater image formation into the rendering pipeline for accurate view synthesis [25, 45]. This helps compensate for the inherent attenuation and optical artifacts in underwater imagery and ensures geometrically consistency. Various image recognition models and 3D reconstruction pipelines also use them for improved visual perception [47, 43].
@@
Note that the image style transfer literature [9, 50] offers solutions that learn artistic or photometric features from data. These methods rely on generative or adversarial networks to adapt one image domain’s style to another [15, 31]. Existing learning pipelines focus on perceptual photorealism [10] or artistic blending [8, 11, 12]), and thus cannot guarantee geometric preservation and depth integrity. While these models are computationally heavy themselves, running physics-based water removal methods [3] in the backend would make it infeasible for end-to-end and/or real-time use [22]. Hence, data augmentation by such domain transfer is not suitable if physical realism and object geometry preservation are crucial, as in 3D view synthesis, scene reconstructions, and robotic mapping tasks [45, 43].
@@
In addition to the theoretical analysis, we present the engineering constructs to tackle the practicalities of waterbody fusion across multiple scenes. We demonstrate how AquaFuse can be used for waterbody crossover, image enhancement by fusing with a clear reference scene, and in a generative pipeline for data augmentation. For qualitative and quantitative validation, we collect data at multiple ocean sites over diverse waterbody types and depth levels ($10^{\prime}$-$80^{\prime}$). Our analyses of $N$ images across six water body types show that AquaFused images preserve over $94\%$ depth consistency and $90$-$95\%$ structural similarity of input scenes. The performance margins vary $1$-$5$% with increasing water depths from $10^{\prime}$ to $80^{\prime}$. We compare the performance with SOTA methods for traditional style transfer, including ASTMAN [13], IEContraAST [8], MCCNet [11], PhotoWCT2 [10] and StyTr2 [12]. With qualitative and quantitative analyses, we demonstrate where traditional learning pipelines fall short for waterbody fusion.
Moreover, we integrate AquaFuse within a 3D view synthesis pipeline using Gaussian Splatting [24] for underwater scene reconstruction. We find that AquaFuse can generate highly detailed 3D reconstructions that represent the original scene’s geometric characteristics in the fused scenes. Experimental results show that AquaFuse (i) performs realistic and accurate waterbody fusion, validated by PSNR (peak signal-to-noise ratio), SSIM (structural similarity index measure), and LPIPS (learned perceptual image patch similarity) scores; (ii) ensures geometric consistency across underwater scenes, validated by SND (surface normal deviation); and (iii) is invariant to dataset bias and overfitting unlike learning-based approaches.
As annotated underwater image databases are scarce, physics-guided data augmentation by AquaFuse can facilitate a multi-fold increase in training samples, which can be used for both 2D image recognition and 3D scene reconstruction [45, 49, 25]. This covers a wide spectrum of robot vision tasks, empowering data-driven models to be more robust and generalizable [48, 1]. Moreover, being a closed-form solution, AquaFuse offers over $342$ FPS ($20$ FPS) inference for $256\times 256$ ($1920\times 1080$) resolutions on a Core™i7 CPU. In fact, it runs on a Raspberry Pi-4 device at over $22$  FPS ($256\times 256$ resolution), demonstrating its utility for real-time active vision on embedded platforms, which no existing style transfer methods can offer.
@@ 2 Background & Preliminaries
Here, $\mathbf{U}$ is the observed underwater image and $\mathbf{I}$ is the clear latent image we want to recover; $\mathbf{\lambda}$ denotes the different wavelengths for RGB channels and $d(\mathbf{x})$ is the actual distance of point $\mathbf{x}$ from camera along the line-of-sight, approximated by depth estimators ($\mathbf{z}$). The illumination map of the medium is an exponential of the form ${\mathbf{E}_{\mathbf{\lambda}}}\sim t_{\mathbf{\lambda}}(\mathbf{x})=e^{-\beta_{\lambda}\cdot d(\mathbf{x})}$, with $\beta_{\mathbf{\lambda}}$ as the uniform attenuation coefficient. Lastly, $\mathbf{B}_{\mathbf{\lambda}}^{\infty}$ represents the background (veiling) light.
The revised underwater image formation model [2, 3] shows that the effects of light absorption can be comparable to light scattering or even dominate it depending on the waterbody. The attenuation coefficient relies not only on the optical properties of water, but also on the sensor, scene irradiance, and imaging range. More importantly, attenuation coefficients of backscattering and direct signal are different: $\beta_{\mathbf{\lambda}}^{B}=-\ln(1-{\mathbf{B}_{\mathbf{\lambda}}}/{\mathbf{B}_{\mathbf{\lambda}}^{\infty}})/\mathbf{z}$ and $\beta_{\mathbf{\lambda}}^{D}=-\ln({(\mathbf{U}_{\mathbf{\lambda}}-\mathbf{B}_{\mathbf{\lambda}})}/{\mathbf{I}_{\mathbf{\lambda}}})/\mathbf{z}$, respectively. The revised model is then expressed by:
@@
where $\mathbf{v}_{D}=\{z,\rho,{E},S_{\lambda},\beta\}$ is the coefficient dependency of direct signal and $\mathbf{v}_{B}=\{{E},S_{\lambda},b,\beta\}$ is coefficient dependency of the backscattered signal. Here, $\rho$ is the reflectance, ${E}$ is the spectrum of ambient light, $S_{\lambda}$ is spectral response, $b$ is beam scattering coefficient, and $\beta$ is beam attenuation coefficient.
In recent years, this revised image formation model has been successfully used for rendering underwater scenes by methods such as SeaThru-NeRF [25], WaterSplatting [30], and SeaSplat [45]. These methods leverage the remarkable advances in 3D radiance fields, specifically NeRF [35] and Gaussian splatting [24]-based rendering. However, drawbacks remain in accounting for multiple scattering or inconsistent/artificial illumination. They also require camera poses to be extracted beforehand, which is often not feasible in practice. Thus, collecting and synthesizing comprehensive data for 3D underwater scene reconstruction research is rather challenging.
@@ 3 AquaFuse Methodology
Figure 2: An illustrative example of our image restoration pipeline; (a) input signal $\mathbf{U}$; (b) ground truth depth; (c) estimated depth $\mathbf{Z}$; (d,e) estimated illumination map for $R$ and $M$ channels; and (f) the restored signal $\mathbf{I}$.
@@ 3.1 Real-time Depth Estimation
A sample example of estimated $\mathbf{Z}\sim\tilde{d}(\cdot)$ is shown in Fig. 2. In our implementation, an additional median blur is applied to reduce noise and preserve edges, which ensures the robustness of background light estimation (see Sec. 3.2). Specifically, we perform $\mathbf{z}=MedianBlur(\mathbf{z},7)\in\mathbf{Z}$ to filter the estimated depth map with a kernel size of $7$.
@@ 4.3.3 Image Enhancement
In Fig. 8, we further demonstrate the influence of $\theta$, the incident angle between incoming light rays and the surface normal. While an exact estimation of $\theta$ is challenging outside controlled environments, this relationship is generally represented using a scalar cosine factor $\cos\theta$ corresponding to the light incidence [37, 39, 41] amount. As Fig. 8 shows, the interpretation for AquaFuse is that $\theta$ can be tuned to control from $\theta\rightarrow 0$ (maximum fusion) to $\theta\rightarrow 90^{\circ}$ (minimum fusion) of waterbody styles from reference image to the content image. Our empirical tests show that $\theta\in[30^{\circ},60^{\circ}]$ generates consistent fusion for most natural underwater scenes.
Figure 8: Effects of incident angle $\theta$ on AquaFuse are shown (with reference $\mathbf{U}7$); lower values of $\theta$ cause more reddish tones (from reference image), while higher values reduce the fusion effects.
@@ 4.3.4 Geometry-preserving Scene Transformation
These visual inspections corroborate our quantitative analyses of the reconstructed scenes. As shown in Table 2, the PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS) scores of AquaFused images are within $1$-$2$% of original scenes. In fact, AquaFuse slightly improves 3D reconstruction quality with a higher PSNR and lower LPIPS at shallow depths ($10^{\prime}$-$30^{\prime}$). For deep-water scenes (over $60^{\prime}$), the scores are very similar to using the original images’ reconstruction quality, which validates our intuition and design objectives. Additionally, 3D reconstructions with AquaFused images show minimal geometric deviations, measured based on the Surface Normal Deviation (SND) metric. As shown in Table 2, the SND is less than $0.3\%$ compared to the original scenes, validating the geometric accuracy of AquaFuse.
Table 2: 3D reconstruction (Gaussian Splatting) performance of AquaFused images versus original images at varied water depths, based on PSNR ($\uparrow$), SSIM ($\uparrow$), and SND ($\downarrow$).
[…]
$10^{\prime}$ Original $22.83$ $0.8104$ $0.2289$ $0.29$
@@
Computational performance analysis. AquaFuse offers a closed-form solution requiring only standard OpenCV libraries, thus offering fast end-to-end inference. As Table 4 indicates, AquaFuse offers $342.3$ FPS ($20.69$ FPS) inference rates for $256\times 256$ ($1920\times 1080$) image resolutions on a basic CPU – which is several orders of magnitude higher than SOTA models. Even on a resource-constrained Raspberry Pi-4 device, AquaFuse offers $22.58$ FPS inference ($44.3$ milliseconds/image) – which no data-driven style transfer methods can offer. These results validate AquaFuse’s feasibility for real-time use on marine robots and standalone sensor nodes.
Table 3: An extension of Table 1, for quantitatively comparing AquaFuse with SOTA style transfer methods. The SSIM and RMSE scores are provided as $mean$$\pm$$variance$.
Depth ASTMAN [13] IEContraAST [8] MCCNet [11]
SSIM ($\uparrow$) RMSE ($\downarrow$) SSIM ($\uparrow$) RMSE ($\downarrow$) SSIM ($\uparrow$) RMSE ($\downarrow$)
$10^{\prime}$ $0.28\pm 0.02$ $51.61\pm 4.69$ $0.59\pm 0.06$ $32.11\pm 5.09$ $0.54\pm 0.06$ $28.63\pm 5.74$
$20^{\prime}$ $0.35\pm 0.02$ $33.02\pm 0.72$ $0.64\pm 0.04$ $24.61\pm 1.29$ $0.63\pm 0.02$ $22.05\pm 1.71$
$30^{\prime}$ $0.21\pm 0.02$ $54.08\pm 2.22$ $0.62\pm 0.03$ $23.09\pm 2.17$ $0.57\pm 0.01$ $20.81\pm 1.99$
$50^{\prime}$ $0.16\pm 0.01$ $51.23\pm 3.77$ $0.59\pm 0.05$ $25.86\pm 4.29$ $0.40\pm 0.12$ $29.21\pm 3.87$
$70^{\prime}$ $0.27\pm 0.03$ $59.45\pm 6.46$ $0.69\pm 0.05$ $29.48\pm 4.38$ $0.58\pm 0.14$ $31.23\pm 4.32$
$80^{\prime}$ $0.36\pm 0.02$ $58.82\pm 7.07$ $0.75\pm 0.05$ $34.71\pm 4.66$ $0.65\pm 0.08$ $32.21\pm 5.03$
Depth StyTr2 [12] PhotoWCT2 [10] AquaFuse
SSIM ($\uparrow$) RMSE ($\downarrow$) SSIM ($\uparrow$) RMSE ($\downarrow$) SSIM ($\uparrow$) RMSE ($\downarrow$)
$10^{\prime}$ $0.30\pm 0.02$ $43.29\pm 1.93$ $0.82\pm 0.07$ $16.16\pm 1.48$ $0.93\pm 0.01$ $5.00\pm 3.66$
$20^{\prime}$ $0.36\pm 0.03$ $27.37\pm 0.44$ $0.88\pm 0.04$ $15.74\pm 1.58$ $0.95\pm 0.00$ $4.46\pm 2.21$
$30^{\prime}$ $0.22\pm 0.03$ $42.16\pm 0.39$ $0.79\pm 0.05$ $15.26\pm 1.57$ $0.93\pm 0.00$ $5.25\pm 1.99$
$50^{\prime}$ $0.16\pm 0.02$ $41.91\pm 1.55$ $0.76\pm 0.06$ $14.67\pm 2.19$ $0.90\pm 0.02$ $5.17\pm 3.52$
$70^{\prime}$ $0.31\pm 0.03$ $44.17\pm 3.33$ $0.82\pm 0.06$ $17.79\pm 2.44$ $0.92\pm 0.02$ $5.58\pm 4.67$
$80^{\prime}$ $0.38\pm 0.02$ $45.83\pm 2.65$ $0.87\pm 0.03$ $22.12\pm 1.33$ $0.90\pm 0.03$ $6.60\pm 6.44$
Figure 10: Comparative qualitative analysis of AquaFuse with SOTA style transfer methods. The input image (content: $\mathbf{U}1$) is transformed by style images ($\mathbf{U}3$, $\mathbf{U}5$, $\mathbf{U}7$, $\mathbf{U}8$) at each row. AquaFuse effectively transfers waterbody style while preserving scene structure.
4.4 Comparative Analysis with Traditional Learning-based Style Transfer Methods
We further investigate SOTA data-driven models for learning waterbody style transfer for three classes (greenish, blueish, and clear) shown in Fig. 6. Results reported in Fig. 10 reveal that transformers and CNN-based models (StyTr2 [12], MCCNet [11], ASTMAN [13]) struggle to preserve the structural integrity of the scene. Although the geometry-aware PhotoWCT2 [10] demonstrates better performance, its generated images exhibit saturated hues and pixel distortions. All models were pre-trained on $118$ K MSCOCO-2017 samples, then fine-tuned on an additional $1,742$ underwater images per class based on their recommended training settings. Quantitative analyses performed on depth-wise datasets corroborate these observations; see Table 3. AquaFuse significantly outperforms competing models across all metrics, validating the robustness and accuracy of its physics-based formulation.
Table 4: Inference rates in FPS ($\uparrow$) for six SOTA methods across various image resolutions on an Intel® Core™-i7 CPU with $16$ GB RAM; ‘$\times$’ represents out-of-memory (or below $0.01$ FPS) cases.
Resolution $256\times 256$ $512\times 512$ $1024\times 1024$ $1920\times 1080$
ASTMAN [13] $0.52$ $0.10$$\times$ $\times$
IEContraAST [8] $0.79$ $0.17$ $0.03$$\times$
MCCNet [11] $0.48$ $0.11$ $0.03$$\times$
StyTr2 [12] $0.12$$\times$ $\times$ $\times$
PhotoWCT2 [10] $0.75$ $0.21$ $0.04$ $0.03$
Sea-thru [3] $1.22$ $0.41$ $0.11$ $0.05$
AquaFuse $342.30$ $84.37$ $20.69$ $9.53$